OpenStudy (anonymous):

Solve the limit: (2(x+h)-2x)/h as h->0

OpenStudy (anonymous):

Are you saying h is 0 and the equation is equal to 0?

OpenStudy (anonymous):

I mean to solve the equation as h approaches 0

OpenStudy (anonymous):

Oh okay i see what you mean

OpenStudy (zzr0ck3r):

\[\frac{2(x+h)-2x}{h}\\=\frac{2x+2h-2x}{h}\\=\frac{2h}{h}=2\] the limit is 2

OpenStudy (anonymous):

Thank you :)